# From Informal to Formal – Incorporating and Evaluating LLMs on Natural Language Requirements to Verifiable Formal Proofs

### v1

@@ Abstract
The research in AI-based formal mathematical reasoning has shown an unstoppable growth trend. These studies have excelled in mathematical competitions like IMO, showing significant progress. However, these studies intertwined multiple skills simultaneously—problem-solving, reasoning, and writing formal specifications—making it hard to precisely identify the LLMs’ strengths and weaknesses in each task. This paper focuses on formal verification, an immediate application scenario of formal reasoning, and decomposes it into six sub-tasks. We constructed 18k high-quality instruction-response pairs across five mainstream formal specification languages (Coq, Lean4, Dafny, ACSL, and TLA+) in six formal-verification-related tasks by distilling gpt-4o. They are split into a 14k+ fine-tuning dataset fm-alpaca and a 4k benchmark fm-bench. We found that LLMs are good at writing proof segments when given either the code, or the detailed description of proof steps. Also, the fine-tuning brought about a nearly threefold improvement at most. And interestingly, we observed that fine-tuning with formal data also enhances abilities in mathematics, reasoning, and coding. We hope our findings inspire further research. Fine-tuned models are released to facilitate subsequent studies.
@@
Jialun Cao1, Yaojie Lu2, Meiziniu Li1, Haoyang Ma1, Haokun Li3, Mengda He3, Cheng Wen4, Le Sun2, Hongyu Zhang5, Shengchao Qin4,6, Shing-Chi Cheung1, Cong Tian6 1 The Hong Kong University of Science and Technology 2 Institute of Software, Chinese Academy of Sciences 3 Fermat Labs, Huawei 4 Guangzhou Institute of Technology, Xidian University 5 Chongqing University 6 ICTT and ISN Laboratory, Xidian University jcaoap@cse.ust.hk, luyaojie@iscas.ac.cn
@@ 1 Introduction
Figure 2: Six tasks towards Informal to Formal Verification
To facilitate the study, we constructed 18k high-quality instruction-response pairs across five mainstream formal specification languages (i.e., Coq, Lean4, Dafny, ACSL, and TLA+) in six formal-verification-related tasks by distilling gpt-4o inspired by prior work Wang et al. (2024); Ding et al. (2023); Wang et al. (2022), then split them into 14k instruction fine-tuning data (fm-alpaca) and 4k benchmarking data (fm-bench). In particular, we provide executable contexts for all these formal specifications and automated validation scripts to validate the correctness of the generated formal proofs inspired from the prior work’s artifact preparation Jimenez et al. (2023). Finally, we release a bunch of fine-tuned LLMs based on three base models (i.e., llama3.1-8B, qwen2.5-7B, and deepseek-coder-7B) in the link  [^2] https://huggingface.co/fm-universe.
@@
2 Design
2.1 Task Formulation
@@ 2.2 Formal Specification Language Selection
First, for Interactive theorem provers which are suitable for developing rigorous mathematical proofs, we consider Coq Huet (1986) and Lean4 Moura and Ullrich (2021) because Coq has been extensively used in academia and research for proving mathematical theorems and in formal verification of software for a long history, while Lean4 garnered considerable attention from the mathematical community Wang et al. (2024); Tao et al. (2023); Avigad et al. (2020) recently.
Second, for programming languages with built-in specification, we consider Dafny Leino (2010) and ACSL Baudin et al. (2021); Cuoq et al. (2012) because they seamlessly integrate specifications (e.g., pre-/post-conditions, loop invariants) within the code, ensuring the correctness through embedded assertions and conditions.
Lastly, for model checking Jhala and Majumdar (2009); Clarke (1997), we consider TLA+ Yu et al. (1999) because it is a representative math-based formal language for modeling algorithms and programs, especially the concurrent and distributed systems.
3 Data Construction
@@ 3.1 Data Preparation
The workflow of data preparation for fm-alpaca and fm-bench is illustrated in Figure 3. The workflow begins with the data collection, where formal proofs in the desired formal specification languages and related configurations and dependencies are gathered from open-source repositories in Github. Then, the formal proofs are extracted from the collected repositories. Next, the proofs go through the data quality assurance check by execution, the proofs that cannot be launched and verified successfully are filtered out. The remaining ones are split into segments (e.g., definition of functions, types, conditions).
@@
After pairing the instruction-response for different tasks, we manually designed five paraphrased task instructions and randomly assigned one for each paired data set to increase instruction diversity and avoid overfitting to certain instructions Lu et al. (2022); Feng et al. (2023).
Finally, we split the prepared data for all the tasks and specification languages into an 8:2 ratio, i.e., 80% for fine-tuning, named fm-alpaca, 20% for benchmarking, named fm-bench.
@@ 6 Code2Proof 70 56 14
deepseek-coder-instruct-v1.5 7B 0.00 2.13 6.78 0.00 7.14 4.29 1.73 3.79 5.56 0.00 9.09 2.38 13.89 28.57 0.00 3.03 13.85 33.33 7.14 14.29
Task-wise Average 6.02 8.54 15.70 16.76 42.86
@@
deepseek-coder-v1.5-fma 7B 0.00 2.13 16.95 0.00 21.43 34.29 25.30 31.06 25.00 84.13 36.36 8.33 19.44 28.57 0.00 1.82 7.69 6.67 21.43 35.71
Task-wise Average 12.72 $\uparrow$ 33.64 $\uparrow$ 17.92 $\uparrow$ 9.31 $\downarrow$ 22.92 $\downarrow$
@@
3.3 Validation Mechanism
For the tasks whose outputs are written in formal specification languages, we verify the full formal proofs against the corresponding verifiers, i.e., Coq, Dafny, Lean4 use their own proving environment; formal specification written in TLA+ can be checked by TLC Yu and Kuppe ; C programs with ACSL specifications can be checked by Frama-C Cuoq et al. (2012); Carvalho et al. (2014). In addition, for the proof segments that cannot be verified independently, for each extracted segment, we prepared a proof template with a placeholder during data preparation. Whenever a generated segment is to be verified, we replace the placeholder in the template with the segment and verify the completed formal proof.
@@ 4.1 Experiment Setup
The experiment is designed to explore three research questions (RQs).
Studied LLMs. We selected eight LLMs as baseline models without fine-tuning, including llama3.1-instruct-8B/70B (Meta, 2024), qwen2.5-instruct-7B/72B (Yang et al., 2024a), qwen2.5-coder-instruct-7B/-32B Hui et al. (2024), deepseek-coder-instruct-7B-v1.5 and deepseek-coder-instruct-33B (Guo et al., 2024). Note that we avoid evaluating the GPT-series LLMs by OpenAI because the descriptions in fm-bench were generated by GPT-4o, making the comparison unfair.
@@
Baseline Fine-tuning Datasets: To distinguish whether the capability improvement is simply because more instruction tuning is applied, we also include two commonly used fine-tuning datasets for comparison. We select UltraChat Ding et al. (2023) and Tulu-V3 Lambert et al. (2024) as baseline fine-tuning datasets. In particular, UltraChat is a large-scale dataset of instructional conversations that contains 1.5 million high-quality multi-turn dialogues and covers a wide range of topics and instructions. Tulu-v3 Lambert et al. (2024) embraces new data that is either carefully manually curated for quality or generated from GPT models. It is an enhancement of its previous versions, Tulu-v2 Ivison et al. (2023) and Tulu-v1 Wang et al. (2023), focusing more on core skills of knowledge recall, reasoning, mathematics, coding, instruction following, general chat, and safety.
Experiment Environment. The experiment uses a high-performance computing environment with the latest GPU technology to ensure precise and efficient model training and evaluation. Each fine-tuning experiment was conducted using 32 Nvidia A100-40G GPUs, while inference experiments were performed with a single Nvidia A100-80G GPU utilizing vLLM (Kwon et al., 2023).
4.2 RQ1. Basic Performance across Formal Specification Tasks
Task-wise: LLMs perform the best in generating proof from code (Code2Proof), with an average of 42.86% Pass@1, followed by ProofInfill (16.76%) and ProofComplete (15.70%). In contrast, LLMs fall short in generating both the entire formal proof (6.02%) and the proof segments (8.54%). We analyzed the failures and found that syntax errors account for a large proportion, with 12.15% failures caused by syntax errors. The observation echoes the motivation of prior work Wang et al. (2024) and is reasonable due to the grammar difference between most formal specification languages and other programming languages like Python. Regarding requirement analysis, as shown in the upper part of Table 4, the Bleu scores between the ground-truth description and LLM-generated ones range from 0.24 to 0.55.
LLM-wise: Horizontally, without fine-tuning, qwen2.5-coder-instruct-32B achieved the best average, followed by qwen2.5-instruct-70B and llama-3.1-instruct-70B.
Model Size – Larger LLMs generally perform better than smaller LLMs. For example, llama3.1-8B only achieved 1.43% in generating TLA+ segments, while llama3.1-70B boosts to 22.86% in the same task. However, there are several exceptions worth noticing, especially for ProofInfill and Code2Proof. For example, llama-3.1-8B achieved 50% in ProofInfill (ACSL), yet the performance drops to 21.43% when using the 70B model. Similar observations can be found in qwen2.5-instruct. The decrease in performance is inherently due to the fine-tuning strategy of these instruction models: the model is trained to excel in generation rather than filling in the blanks Fried et al. (2022). Additionally, we conducted a more detailed examination of generation by models of different scales and observed that larger models tend to fill in the proof segments that not only look more plausibly correct and well-organized but also tend to include extra content. The additional content, yet, is either redundant, as it repeats information that appears in the subsequent proof, or is incomplete. This thus leads to situations where smaller models perform correctly in ProofInfill and Code2Proof where larger models do not. Promisingly, recent model developers have noticed such conundrums and refined their fine-tuning strategy for fill-in-the-middle tasks Guo et al. (2024).
@@
Table 5 shows the LLMs’ performance across formal specification languages in the task of generating proof segments (SegGen). This task accounts for the most instructions and serves as basic capability for other proof generation tasks. We can see that LLMs perform the best in ACSL (an average of 34.92%), followed by Dafny (15.97%) while performing unsatisfactorily in other formal specification languages. The observation is reasonable because the syntax of ACSL is basically an annotation of C language, while Dafny shares similar grammar as C# and Java. Thus, generating proof segments in ACSL and Dafny is generally easier than generating other formal specification languages.
In addition, we explore whether increasing the attempts (1 $\rightarrow$ 5) with a higher temperature (0.0 $\rightarrow$ 0.2) and in-context learning (i.e., 3-shot) could bring about improvement. The improvement ratios are shown in red in Table 5. Clear that the results of Pass@5 are better than those of Pass@1, with an average score increase from 10.82% (Dafny) to 63.64% (ACSL) in different languages. Moreover, when using 3-shot learning, the performance increases dramatically, with 51.33% (Dafny) to over five times (ACSL) improvement compared with zero-shot Pass@5. The results indicate the potential of in-context learning in generating correct formal specification languages.
@@
On the other hand, drops can be observed in fill-in-the-middle tasks (i.e., ProofInfill and Code2Proof). The results echo the observation made in RQ1 (Section 4.2), where the large LLMs perform worse than small LLMs in fill-in-the-middle tasks. The results also indicate the necessity of adopting different fine-tuning strategies other than instruction tuning only.
Fine-tuning Datasets: Take a closer look at the LLMs fine-tuned with general-purpose datasets (i.e., llama3.1-ultrachat and llama3.1-tulu) in Table 3, with them only, no or opposite effects can be observed. The results indicate the complementarity of fm-alpaca and existing general-purpose fine-tuning datasets. Additionally, by combining with general-purpose datasets, the performance can be further improved (i.e., llama3.1-ultrachat-fma and llama3.1-tulu-fma).
@@
Finally, we explore whether fine-tuning with fm-alpaca could benefit related capabilities. Table 6 shows the experiment results. In particular, the base model is llama3.1-8B, fine-tuned under two base fine-tuned datasets with and without fm-alpaca. On average, with fm-alpaca, an increase of 1.37% to 5.15% can be observed. Interestingly, a dramatic increase (62.53%) can be observed in HumanEval compared with the performance of the model that is only fine-tuned with UltraChat. Similarly, the performance in Math and Reasoning also witness slight improvements. The experiment also indicates the potential of improving LLMs’ coding, reasoning, and math capabilities by feeding more formal data.

### v3

@@ Abstract
The research in AI-based formal mathematical reasoning has shown an unstoppable growth trend. These studies have excelled in mathematical competitions like IMO and have made significant progress. However, these studies intertwined multiple skills simultaneously—problem-solving, reasoning, and writing formal specifications—making it hard to precisely identify the LLMs’ strengths and weaknesses in each task. This paper focuses on formal verification, an immediate application scenario of formal reasoning, and breaks it down into sub-tasks. We constructed 18k high-quality instruction-response pairs across five mainstream formal specification languages (Coq, Lean4, Dafny, ACSL, and TLA+) in six tasks by distilling gpt-4o and evaluated against ten open-sourced LLMs, including recent popular DeepSeek-R1. We found that LLMs are good at writing proof segments when given either the code, or the detailed description of proof steps. Also, the fine-tuning brought about a nearly threefold improvement at most. Interestingly, we observed that fine-tuning with formal data also enhances mathematics, reasoning, and coding capabilities. Fine-tuned models are released to facilitate subsequent studies at https://huggingface.co/fm-universe.
@@
Jialun Cao1, Yaojie Lu2, Meiziniu Li1, Haoyang Ma1, Haokun Li3, Mengda He3, Cheng Wen3, Le Sun2, Hongyu Zhang4, Shengchao Qin3,5, Shing-Chi Cheung1, Cong Tian5 1 The Hong Kong University of Science and Technology 2 Institute of Software, Chinese Academy of Sciences 3 Guangzhou Institute of Technology, Xidian University 4 Chongqing University 5 ICTT and ISN Laboratory, Xidian University jcaoap@cse.ust.hk, luyaojie@iscas.ac.cn
@@ 1 Introduction
To facilitate the study, we constructed 18k high-quality instruction-response pairs across five formal specification languages (i.e., Coq, Lean4, Dafny, ACSL, and TLA+) in six formal-verification-related tasks by distilling gpt-4o inspired by prior work Wang et al. (2024); Ding et al. (2023); Wang et al. (2022), then split them into 14k instruction fine-tuning data (fm-alpaca) and 4k benchmarking data (fm-bench). In particular, we provide executable contexts for all these formal specifications and automated validation scripts to validate the correctness of the generated formal proofs inspired from the prior work’s artifact preparation Jimenez et al. (2023). Finally, we release the fine-tuned LLMs based on three base models at https://huggingface.co/fm-universe.
@@
2 Task Formulation
Figure 2 illustrates the six sub-tasks. We elaborate on them in detail as follows.
@@ 2 Task Formulation
Figure 2: Six tasks towards Informal to Formal Verification
3 Data Construction
@@ 3.1 Formal Specification Language Selection
First, for interactive theorem provers which are suitable for developing rigorous mathematical proofs, we consider Coq Huet (1986) and Lean4 Moura and Ullrich (2021) because Coq has been extensively used in academia and research for proving mathematical theorems and in formal verification of software for a long history, while Lean4 garnered considerable attention from the mathematical community Wang et al. (2024); Tao et al. (2023); Avigad et al. (2020) recently. Second, for programming languages with built-in specification, we consider Dafny Leino (2010) and ACSL Baudin et al. (2021); Cuoq et al. (2012) because they seamlessly integrate specifications (e.g., pre-/post-conditions, loop invariants) within the code, ensuring the correctness through embedded assertions and conditions. Lastly, for model checking Jhala and Majumdar (2009); Clarke (1997), we consider TLA+ Yu et al. (1999) since it is a representative math-based formal language for modeling algorithms and programs such as concurrent and distributed systems.
@@ 3.2 Data Preparation
The workflow of data preparation for fm-alpaca and fm-bench is illustrated in Figure 3. The workflow begins with the data collection, where formal proofs in the desired formal specification languages and related configurations and dependencies are gathered from open-source repositories in Github. Then, formal proofs are extracted from the collected repositories. Next, the proofs go through the data quality assurance check by execution, the proofs that cannot be verified successfully are filtered out. The remaining ones are split into segments (e.g., definition of functions or conditions).
@@
After pairing the instruction-response for different tasks, we manually designed five task instructions for each task and randomly assigned one for each paired data to increase instruction diversity and avoid overfitting to certain instructions Lu et al. (2022); Feng et al. (2023); Sanh et al. (2022).
@@ 6 Code2Proof 70 56 14
3.4 Validation Mechanism
For the tasks whose outputs are written in formal specification languages, we verify the full proofs against the corresponding verifiers, i.e., Coq, Dafny, Lean4 use their own proving environment; formal specification written in TLA+ can be checked by TLC Yu and Kuppe ; C programs with ACSL specifications can be checked by Frama-C Cuoq et al. (2012); Carvalho et al. (2014). Also, for the proof segments that cannot be verified independently, for each extracted segment, we prepared a proof template with a placeholder during data preparation. Whenever a generated segment is to be verified, we replace the placeholder in the template with the segment and verify the completed formal proof.
@@ 3.4 Validation Mechanism
LLMs Size TLA Coq Lean Dafny ACSL TLA Coq Lean Dafny ACSL TLA Coq Lean Dafny TLA Coq Lean Dafny ACSL ACSL Ave
@@
llama3.1-instruct 8B 0.00 4.26 1.69 0.00 14.29 1.43 1.19 6.06 8.33 0.00 0.00 3.57 5.56 28.57 0.00 4.85 7.69 13.33 50.00 71.43 11.11
llama3.1-instruct 70B 7.69 6.38 10.17 20.00 7.14 22.86 4.86 13.64 11.11 39.68 27.27 3.57 13.89 28.57 0.00 7.27 24.62 13.33 21.43 21.43 15.25
qwen2.5-instruct 7B 0.00 2.13 5.08 0.00 21.43 1.43 1.05 7.58 5.56 1.59 0.00 13.10 19.44 28.57 0.00 5.45 18.46 13.33 50.00 57.14 12.57
qwen2.5-instruct 72B 7.69 6.38 18.64 20.00 21.43 12.86 3.29 12.12 25.00 12.70 0.00 11.90 33.33 28.57 0.00 9.70 24.62 26.67 35.71 57.14 18.39
qwen2.5-coder-instruct 7B 0.00 2.13 6.78 0.00 0.00 2.86 2.14 12.12 11.11 3.17 0.00 4.76 11.11 14.29 0.00 3.64 23.08 20.00 50.00 57.14 11.22
qwen2.5-coder-instruct 32B 0.00 6.38 11.86 0.00 21.43 12.86 4.07 9.85 25.00 26.98 27.27 5.95 36.11 42.86 2.94 7.88 33.85 20.00 71.43 57.14 21.19
starcoder-instruct 15B 7.69 2.13 8.47 30.00 28.57 27.14 1.97 6.06 27.78 0.00 27.27 14.29 30.56 42.86 0.00 11.52 27.69 20.00 35.71 28.57 18.27
deepseek-coder-instruct-v1.5 7B 0.00 2.13 6.78 0.00 7.14 4.29 1.73 3.79 5.56 0.00 9.09 2.38 13.89 28.57 0.00 3.03 13.85 33.33 7.14 14.29 7.85
deepseek-coder-instruct 33B 0.00 0.00 1.69 0.00 0.00 2.86 1.94 4.55 11.11 3.17 0.00 3.57 13.89 42.86 0.00 2.42 23.08 6.67 21.43 7.14 7.32
deepseek-R1 671B 15.38 8.51 25.42 30 35.71 22.86 10.7 21.21 22.22 49.21 45.45 17.86 36.11 28.57 5.88 13.33 40 6.67 42.86 64.29 27.11
Task-wise Average 8.65 10.61 18.44 17.38 43.57
@@
llama3.1-fma 8B 0.00 6.38 8.47 20.00 57.14 41.43 29.98 21.97 25.00 90.48 36.36 8.33 11.11 28.57 0.00 0.00 7.69 0.00 21.43 21.43 21.79
llama3.1-ultrachat 8B 0.00 0.00 1.69 0.00 7.14 0.00 0.00 3.79 0.00 0.00 0.00 5.95 13.89 14.29 0.00 4.85 9.23 0.00 14.29 0.00 3.76
llama3.1-ultrachat-fma 8B 0.00 8.51 10.17 30.00 50.00 41.43 35.72 29.55 33.33 95.24 18.18 8.33 11.11 28.57 2.94 0.00 7.69 0.00 28.57 35.71 23.75
llama3.1-tulu 8B 0.00 0.00 1.69 0.00 7.14 0.00 0.71 6.06 0.00 0.00 0.00 5.95 8.33 28.57 0.00 1.82 10.77 0.00 57.14 50.00 8.91
llama3.1-tulu-fma 8B 0.00 4.26 11.86 30.00 35.71 42.86 36.71 27.27 36.11 98.41 18.18 10.71 11.11 42.86 2.94 2.42 6.15 0.00 35.71 50.00 25.16
qwen2.5-fma 7B 0.00 4.26 11.86 10.00 71.43 38.57 27.95 27.27 22.22 87.30 27.27 9.52 13.89 28.57 0.00 0.00 7.69 0.00 35.71 21.43 22.25
qwen2.5-coder-fma 7B 0.00 6.38 18.64 20.00 35.71 44.29 36.50 34.09 33.33 98.41 36.36 9.52 16.67 28.57 5.88 1.82 13.85 13.33 42.86 50.00 27.31
deepseek-coder-v1.5-fma 7B 0.00 2.13 16.95 0.00 21.43 34.29 25.30 31.06 25.00 84.13 36.36 8.33 19.44 28.57 0.00 1.82 7.69 6.67 21.43 35.71 20.32
Task-wise Average 12.72 $\uparrow$ 33.64 $\uparrow$ 17.92 9.31 22.92
@@ 4.1 Experiment Setup
Studied LLMs. We selected ten LLMs as baselines without fine-tuning, including llama3.1-instruct-8B/70B (Meta, 2024), qwen2.5-instruct-7B/72B (Yang et al., 2024a), qwen2.5-coder-instruct-7B/-32B Hui et al. (2024), starcoder-instruct-15B Lozhkov et al. (2024), deepseek-coder-instruct-7B-v1.5, deepseek-coder-instruct-33B (Guo et al., 2024), and deepseek-R1 Guo et al. (2025). Note that we avoid evaluating the GPT-series LLMs by OpenAI because the descriptions in fm-bench were generated by GPT-4o, making the evaluation fairer.
@@
Baseline Fine-tuning Datasets: To distinguish whether the capability improvement is simply because more instruction tuning is applied, we also include two commonly used fine-tuning datasets for comparison. We select UltraChat Ding et al. (2023) and Tulu-V3 Lambert et al. (2024) as baseline fine-tuning datasets for their popularity. In particular, UltraChat is a large-scale dataset of instructional conversations that contains 1.5 million high-quality multi-turn dialogues and covers a wide range of topics and instructions. Tulu-v3 Lambert et al. (2024) embraces new data that is either carefully manually curated for quality or generated from GPT models. It is an enhancement of its previous versions Ivison et al. (2023); Wang et al. (2023), focusing more on core skills of knowledge recall, reasoning, mathematics, coding, instruction following, general chat, and safety.
Experiment Environment. The fine-tuning experiment was conducted on 32 Nvidia A100-40G GPUs, while inference was on a single Nvidia A100-80G GPU with vLLM (Kwon et al., 2023).
4.2 RQ1. Basic Performance across Formal Specification Tasks
Task-wise: LLMs perform the best in generating proof from code (Code2Proof), with an average of 43.57% Pass@1, followed by ProofComplete (18.44%) and ProofInfill (17.38%). In contrast, LLMs fall short in generating both the entire formal proof (8.65%) and the proof segments (10.61%). We analyzed the failures and found that syntax errors account for a large proportion, with 12.15% failures caused by syntax errors (Appendix B). The observation echoes the motivation of prior work Wang et al. (2024) and is reasonable due to the grammar difference between most formal specification languages and other programming languages like Python. Regarding requirement analysis, as shown in the upper part of Table 4, the Bleu scores between the ground-truth description and LLM-generated ones range from 0.24 to 0.55. LLM-wise: Without fine-tuning, DeepSeek-R1 achieved the best average (27.11%), followed by qwen2.5-coder-instruct-32B (21.19%).
Model Size – Larger LLMs generally perform better than smaller LLMs. For example, llama3.1-8B only achieved 1.43% in generating TLA+ segments, while llama3.1-70B boosts to 22.86% in the same task. However, there are several exceptions worth noticing, especially for ProofInfill and Code2Proof. For example, llama-3.1-8B achieved 50% in ProofInfill (ACSL), yet the performance drops to 21.43% using the 70B model. Similar observations can be found in qwen2.5-instruct. The decrease in performance is inherently due to the fine-tuning strategy of these instruction models: they are trained to excel in generating rather than filling in the blanks Fried et al. (2022). Also, we conducted a more detailed examination of generated segments and observed that larger LLMs tend to fill in the proof segments that not only look more plausibly correct and well-organized but also include extra content. The additional content, yet, is either redundant, as it repeats information that appears in the subsequent proof, or is incomplete. Promisingly, recent model developers have noticed such conundrums and refined their fine-tuning strategy for fill-in-the-middle tasks Guo et al. (2024).
@@
Table 5 shows the LLMs’ performance across formal specification languages in the task of generating proof segments (SegGen). This task accounts for the most instructions and serves as the basic capability for other proof generation tasks. We can see that LLMs perform the best in ACSL (average: 34.92%), followed by Dafny (15.97%) while performing unsatisfactorily in other formal specification languages. The observation is reasonable because the syntax of ACSL is basically an annotation of C language, while Dafny shares similar grammar as C# and Java. Thus, generating proof segments in ACSL and Dafny is generally easier than generating other specification languages.
In addition, we explore whether increasing the attempts (1 $\rightarrow$ 5) with a higher temperature (0.0 $\rightarrow$ 0.2) and in-context learning could bring about improvement. The improvement ratios are shown in red in Table 5. The results of Pass@5 are better than those of Pass@1, with an average score increase from 10.82% (Dafny) to 63.64% (ACSL) in different languages. Moreover, when using 3-shot, the performance increases dramatically, with 51.33% (Dafny) to over five times (ACSL) improvement compared with zero-shot Pass@5. The results indicate the potential of in-context learning in generating correct specification languages.
@@
Yet, drops can be observed in fill-in-the-middle tasks (i.e., ProofInfill and Code2Proof). The results echo the observation made in RQ1 (Section 4.2), where the large LLMs perform worse than small LLMs in fill-in-the-middle tasks. The results also indicate the necessity of adopting different fine-tuning strategies other than instruction tuning only.
Fine-tuning Datasets: Take a closer look at the LLMs fine-tuned with general-purpose datasets (i.e., llama3.1-ultrachat and llama3.1-tulu) in Table 3, with them only, no or opposite effects can be observed. The results indicate the complementarity of fm-alpaca and existing general-purpose fine-tuning datasets. Additionally, by combining with general-purpose datasets, the performance can be further improved (e.g., llama3.1-tulu-fma).
@@
Finally, we explore whether fine-tuning with fm-alpaca could benefit related capabilities. Table 6 shows the results. The base model is llama3.1-8B, fine-tuned under two base fine-tuned datasets with and without fm-alpaca. On average, with fm-alpaca, an increase of 1.37% to 5.15% can be observed. Interestingly, a dramatic increase (62.53%) can be observed in HumanEval compared with the performance of the model that is only fine-tuned with UltraChat. The experiment may indicate that feeding more formal data may improve LLMs’ coding, reasoning, and math capabilities.